Given the equation of a circle x
2
+ y2
– 14x – 16y + 88 = 0. Determine the following
a. Center of the circle
b. Radius
c. Area
d. Perimeter
1
Expert's answer
2020-05-06T18:23:53-0400

The given equation of circle is "x^2 + y^2 \u2013 14x \u2013 16y + 88 = 0" . Let us rewrite it in a form "(x-x_0)^2 + (y-y_0)^2 = R^2, \quad (1)"

where "x_0" and "y_0" are the coordinates of the center of the circle.

If we expand the brackets in (1) and subtract "R^2" from the both sides, we'll get

"x^2-2xx_0 + x_0^2 + y^2 - 2yy_0 + y_0^2 - R^2 = 0."

Next we compare this formula and the given equation. We can see that "-2xx_0 = -14x, \; -2yy_0 = -16y,"

so "x_0 = 7, \;\; y_0 = 8."

Therefore, "x_0^2+y_0^2 = 113." From (1) and the given equation we get

"x_0^2+y_0^2 - R^2 = 88".

Therefore, "113-R^2 = 88, \;\; R^2 = 25, \;\; R=5."

a) Center of the circle is (7, 8)

b) Radius is 5.

c) The area of the circle can be calculated as "S=\pi R^2 = \pi\cdot 5^2 = 25\pi \approx 78.54."

d) Perimeter of the circle can be calculated as "P=2\pi R = 2\pi \cdot5 = 10\pi \approx 31.42."